A uniform sphere of radius 0.5m and mass 15kg rotate with the xy plane about axis through the center.find the angular momentum when the angular velocity is 3rads/ seconds


1
Expert's answer
2021-08-04T04:40:44-0400

"w =3\frac{rad}{s}"

"r =0.5m"

"m = 15kg"

"L = I_zw"

"Iz=\frac{2}{5}mr^2= \frac{2}{5}15*0.5^2= 1.5"

"L = 1.5*3 = 4.5"

Answer : L = 4.5